OpenStudy (anonymous):

Use the compound interest formula to compute the total amount accumulated. $10,000 for 20 years at 5% compounded annually A. $20,000.00 B. $25,269.50 C. $19,500.00 D. $26,532.98

OpenStudy (anonymous):

it doesnt say

OpenStudy (phi):

The equation is FV= PV*(1+i)^n where FV is the future value PV is the present value i is the annual interest (as a decimal) n is the number of years. Can you replace all the letters with the numbers given in the problem?

OpenStudy (anonymous):

can you just tell me if it abcd? please

OpenStudy (anonymous):

No @kaitlyncosta - that's not what we're here for. @phi and @Chlorophyll have given you enough to start. They (and we) are not here to simply tell you abcd, sorry.

OpenStudy (anonymous):

I don't get it? Would it be C. $19,500.00 ?

OpenStudy (phi):

@brittkay12 It is more important to understand how to do it than just write down an answer. How you do it, is first match numbers to the variables in the formula so what are the numbers (from the problem) for each of these? PV is the present value i is the annual interest (as a decimal) n is the number of years.

OpenStudy (anonymous):

Wait!! I think I got it.. It's D.!!?!

OpenStudy (phi):

your numbers look good. next step is put them into the formula FV= PV*(1+i)^n FV= 10000*(1+0.05)^20 1+0.05 is 1.05 FV= 10000*1.05^20 can you get the answer using a calculator?

OpenStudy (anonymous):

the answer i got was d ... i did 1.05^20 then multiplied by 10000

OpenStudy (anonymous):

thanks @phi